OpenStudy (anonymous):

Medal and a fan for helping! 3x+y=-13 -2x+5y=-54

OpenStudy (anonymous):

y= -13-3x -2x +5(-13-3x)=-54 -2x +65-15x=-54 -17x=-54-65 -17x=-119 x=119/17 y=-13-(3*119/17) i don't have a calculator in here hope haVE helped

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

did I HELP? ask anything you want if you want help

OpenStudy (anonymous):

Yes! Can you help me with a few more?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

what do you want help for

OpenStudy (anonymous):

2. 5x+4y=-83 3x-3y=-12

OpenStudy (anonymous):

3x=-12+3y x=(3y-12)/3 x=y-4 5(y-4)+4y=-83 5y-20 +4y=-83 9y=-83+20 9y=-63 y=-63/9 x=-63/9-4 what degree are you?